OpenStudy (anonymous):

solve for d when 2x-2yd=y+xd

jimthompson5910 (jim_thompson5910):

2x-2yd=y+xd 2x-y=xd+2yd 2x-y=d(x+2y) (2x-y)/(x+2y)=d So the answer is \[\large d=\frac{2x-y}{x+2y}\]